# Integrated Sensing and Edge AI: Realizing Intelligent Perception in 6G

### v1

@@ IV-A The Design Principles
ISEA aims to integrate sensing, AI, and communications to achieve optimized task performance. The general principle of ISEA design is to process and transmit task-relevant sensory information for various applications with improved accuracy, latency, and energy efficiency. It aligns with not only task-oriented communications of 6G but also the design principle of E2E perception applications. In the sequel, we break down the general principle into several key aspects that distinguish ISEA from conventional mobile communication paradigms, such as semantic communications and ISAC.
@@ V-A1 Key Components
V-A2 Computing Pipeline and AI Backbones
While ISEA is envisioned to support various sensing tasks using different AI models, they share similarities in the computing stage. A typical computing pipeline in the ISEA system includes data pre-processing, feature extraction, feature selection and compression, sensory fusion, and task-specific inference.
@@ V-A2 Computing Pipeline and AI Backbones
Depending on the heterogeneous computing capabilities and communication links, stages in this pipeline can be deployed onto different components in ISEA, resulting in various paradigms as detailed in Section V-B.
@@ V-B5 Hierarchical ISEA
TABLE I: Summary of related works on digital air interface for ISEA
@@
[127] An collaborative ISEA system where an access point (AP) aggregates data from multiple sensors and broadcasts the fused data Total execution time Information extractor design and a task-oriented channel allocation mechanism
@@ VI-C1 Power Control and Bandwidth Allocation
The allocation of power and bandwidth resources to a given set of sensors shall consider the heterogeneous channels, data qualities, and computation capabilities of different sensors to meet the ISEA task performance requirement. In [125], an edge AI-assisted autonomous driving system is considered, where connected vehicles are scheduled to upload sensing data to multiple servers for perception model training. To maximize a defined metric, i.e., quality of training, the power and bandwidth allocation to different vehicles and time slots are jointly optimized. In [126], the authors consider ISAC settings where a BS simultaneously senses and transmits data packets, arriving in a Poisson process, to multiple users with a multi-functional OFDM signal. The subcarrier and power allocation problem is formulated to minimize the long-term average power consumption subject to sensing performance and queueing stability conditions and is solved via Lyapunov optimization. In a collaborative ISEA system with a fusion center considered in[127], the inadequacy of uplink communication resources for uploading all sensing information necessitates information extraction, agent selection, and subchannel allocation according to channel states and importance of messages. The information extraction module is trained using the information bottleneck principle, the subchannel allocation is optimized to maximize the expected information gain, and message fusion is designed with a double-attention mechanism. The authors of [128] first propose a primal-descent dual ascent training algorithm for FL under hybrid data partition, which involves several groups of sensors with different types, to resolve the data dependence issue. Then, the optimization problem of sensor scheduling, bandwidth allocation, and quantization is formulated to minimize the training residual term and resource consumption, which is solved in a decentralized manner.
@@ VI-C3 Joint Computation-Communication Resource Allocation
In ISEA systems, computation and communication resources can both be bottlenecks of task performance, especially in view of time-varying channel conditions and limited on-sensor computation capability. This requires a joint controller coordinating communication and computation operations to ensure full resource utilization. In [130], a point-to-point split inference system involving a UAV and a BS is considered, where the UAV can switch between two modes, i.e., uploading raw data or extracted features, with different C2 resource consumption. Such mode switching is jointly optimized with the transmit power allocation for multiple timeslots and computational speed to minimize the total energy consumption. The resultant optimization problem is solved by embedding the optimal power allocation into a tiny reinforcement learning algorithm for optimal mode selection. In light of the tradeoff between reducing communication cost and preventing semantic loss as regulated by the compression ratio, in [131] a joint ratio selection and resource allocation problem is formulated for a multi-user semantic communication system aiming at supporting inference tasks for users. The solution approach is through alternative minimization that iterates between compression ratios and resource allocation. Further, a user selection algorithm is designed via branch and bound under communication resource constraints, considering that local tasks have different priority levels. Considering split ISEA where multi-user performs inference offloading with heterogeneous accuracy and latency requirements, the authors in [56] propose integrated batching and early exiting to maximize the system inference throughput. Under a total bandwidth constraint and different channel states, the optimized allocation of computation and communication resources for inference throughput maximization is realized via a low-complexity sub-optimal algorithm and a tree search-based optimal algorithm. For collaborative ISEA, the splitting point between sensor and server sub-models can be jointly optimized with C2 resources for minimized energy consumption subject to delay constraints[132].
@@ VII AirComp-based Air Interface for Integrated Sensing and Edge AI
TABLE II: Summary of Related Works on AirComp-based Air Interface for ISEA
@@ VII-C1 Robust AirComp
In addition, AirComp performance is prone to interference from jammers or neighboring cells, which may lead to sudden drops in the sensing task performance, i.e., AI outage. A set of RRM techniques has been developed for interference management. In[153], a scheme called simultaneous signal-and-interference alignment is proposed to manage inference in a two-cell system. It divides the MIMO channel space into an interference alignment space and signal alignment space, and by constraining the interference into the former, the number of interference-free dimensions for AirComp can be maximized. In [135], the centralized power control problem for a multi-cell AirComp network is tackled by profiling the Pareto front of multi-cell MSEs and solving a sequence of second-order cone programming problems, and a distributed algorithm designed by exploiting the interference temperature technique. For multicluster networks, [154] developed a centralized branch-and-bound algorithm and a distributed algorithm for the transceiver optimization for transceiver design to minimize the aggregation MSE accounting for both signal misalignment error and interference. A novel technique, termed spectrum breathing, is developed in [155], which features cascaded gradient pruning and spread spectrum to reduce interference under a given bandwidth budget.
@@ VIII Advanced Signal Processing for Integrated Sensing and Edge AI
TABLE III: Summary of Related Works on Advanced Signal Processing for ISEA
@@ VIII-B ISEA with Multi-Functional Waveforms
From a physical layer perspective, advancing ISEA critically relies on the foundational design of waveforms. Multi-functional waveforms are sophisticated signal designs that integrate multiple functions, such as sensing, communication, localization, and radar, into a unified structure, leveraging advanced signal processing to optimize performance. Previous attempts in this field focused on extending the functionalities of existing systems. For example, communication symbols were modulated onto traditional radar waveforms, such as space-time coding over orthogonal radar waveforms, code shift keying using optimized coded radar signals, and embedding communication symbols in frequency-hopping multi-input multi-output (FH-MIMO) radar systems [187, 188, 189]. Some early works in this field explored the performance of sensing using signals generated through linear frequency modulation (LFM) and continuous phase modulation [190, 191]. These schemes aimed to utilize existing architectures to meet the limited demands of additional functionalities. More recent research focuses on the novel design of multi-functional waveforms for simultaneous monostatic sensing and communication (SMSC). In SMSC, a transceiver array emits dual-functional signals to illuminate the scene and gather environmental information through echo signals. At the same time, forward signals are reflected to a communication receiver for information delivery. The advantage of SMSC is that it allows the uncertainty of communication messages to be incorporated into the transmitted waveforms without compromising the coherent processing of monostatic sensing. This eliminates the need for cooperation overhead [192]. However, the distinct requirements of sensing and communication pose challenges in the design of multi-function waveforms, such as the Cramér-Rao bound (CRB) and data rates, leading to a tradeoff in performance [192, 209]. To address these tradeoffs, research has focused on beamforming techniques, where beam patterns are formed in a joint radar-communication MIMO system based on the directions of sensing targets and communication users to mitigate mutual interference [15, 194, 195, 196]. Additionally, time-frequency waveforms have been revisited to optimize subcarrier powers in a specific time-frequency region of interest, achieving a favorable performance tradeoff between radar and communications [197].
@@ IX-A3 Reconfigurable Intelligent Surfaces
  • Joint beamforming techniques: Reflective control of the RIS must be jointly designed with conventional transmit beamforming to maximize sensing performance, rather than simply focusing on maximizing SNR as in typical communication scenarios. For instance, [231] demonstrates how adaptive RIS beamforming can maximize the received signal power at RIS sensors—a crucial metric for sensing. In general, beamforming strategies should be explored to balance communication efficiency and sensing accuracy.
@@ X-C Ultra-Low-Latency ISEA
However, existing SPT approaches targeting URLLC fall short of meeting the ultra-LoLa ISEA requirements. This shortcoming arises because ISEA applications are typically data-intensive and often require the transmission of high-dimensional features, which conflicts with SPT’s limited payload sizes and leads to communication bottlenecks. The problem is further exacerbated for tasks subject to strict latency requirements. Moreover, the E2E performance of sensing tasks cannot be accurately measured through decoding error probability alone, as that metric is more suited to assessing communication reliability. To address these challenges, the first framework that enables ultra-LoLa ISEA leverages the interplay between SPT and multi-view sensing—improving accuracy while adhering to stringent latency constraints [258]. Within such a framework, a fundamental tradeoff is identified and optimized between communication reliability and the number of sensing views, both of which are governed by the packet length. This integration of SPT and ISEA opens numerous opportunities for future work. One key direction is the simultaneous improvement of accuracy and reduction of latency through a joint optimization of packet length and coding rate, which requires further exploration into the effects of feature dimensionality and quantization on packet selection. From an energy-efficiency standpoint, designing power control techniques that accommodate both fading channels and feature importance is also a promising avenue. Additionally, extending the ultra-LoLa inference paradigm to incorporate AirComp constitutes another compelling research direction.

### v2

@@ III-D3 Datasets
III-E Lessons Learned
  • A wide variety of emerging 6G applications and techniques requires intelligent perception for environment awareness, semantic understanding, and rapid response. Their real-time and mission-critical nature results in unprecedented QoS requirements, including millisecond-level communication-and-computation latency, ultra-high reliability in E2E task performance, and scalability to thousands of cooperating devices[105].
  • These requirements cannot be satisfied with existing 5G networks, which focuses on creating high-throughput bit pipes but are suboptimal in E2E task metrics.
  • ISEA, on the contrary, provides a versatile framework for intelligent perception provisioning in edge networks with E2E optimality. Standardization organization and industry leaders have recognized the potential of ISEA and initiated projects and prototypes that pave the way for real-world implementations of ISEA[97, 98]. Furthermore, a variety of datasets [99, 100, 101, 102, 103, 104] have been made available to simulate multi-modal sensing data acquisition, mobility, and wireless propagation in real-world environments, facilitating benchmarking of ISEA performance. These insights indicate that ISEA represents a promising research direction in the context of 6G technologies.
@@ IV-A The Design Principles
ISEA aims to integrate sensing, AI, and communications to achieve optimized task performance. The general principle of ISEA design is to process and transmit task-relevant sensory information for various applications with improved accuracy, latency, and energy efficiency. It aligns with not only task-oriented communications of 6G but also the design principle of E2E perception applications. In the sequel, we break down the general principle into several key aspects that distinguish ISEA from conventional rate-oriented mobile communication systems and other emerging paradigms such as semantic communications and ISAC.
@@ IV-B4 Tradeoffs
IV-C Lessons Learned
  • Different from the rate-maximization principle in conventional mobile networks, ISEA aims at optimized E2E task performance, e.g., sensing accuracy, E2E latency, and energy efficiency. This involves the coordinated integration of sensing, AI, and communication to efficiently process and transmit task-relevant sensory data.
  • Specifically, communication links are designed to be aware of the semantics and characteristics of the conveyed sensing data. By leveraging the interaction between these factors and the time-varying network connectivity, ISEA implements modality-aware communication schemes that outperforms modality-agnostic approaches in efficiency. Second, ISEA adopts an integrated communication and computing design to optimize resource allocation and develop transmission protocols tailored for AI-empowered sensing models. Finally, all ISEA designs and optimizations target the E2E sensing task performance, allowing for the utilization of challenges in rate-oriented communication systems, such as channel noise and interference, to benefit downstream tasks.
  • Following the task-oriented principle, the ISEA performance is evaluated with a broad set of metrics that go beyond the scope of existing communication systems. Notably, sensing metrics play a crucial role in ISEA design, encompassing accuracy/uncertainty, estimation error, coverage, resolution, and timeliness. Communication metrics, such as throughput, latency, energy efficiency, and mobility, are also integral to ISEA but are adapted to align more closely with downstream tasks. For instance, throughput in ISEA extends beyond mere transmission efficiency to encompass rates of sensing semantic transmission[107] and task completion[62]. Additionally, computation metrics like time complexity and space complexity must be factored in for end-to-end system functionality. While achieving high performance across all metrics may not always be achievable, ISEA designers must strike a balance among these metrics within resource constraints to meet the needs of specific tasks.
@@ V-A1 Key Components
V-A2 Computation Operations and AI Backbones
While ISEA is envisioned to support various sensing tasks using different AI models, they share similarities in the computing stage. A typical pipeline of computation operations in the ISEA system includes data pre-processing, feature extraction, feature selection and compression, sensory fusion, and task-specific inference.
@@ V-A2 Computation Operations and AI Backbones
Depending on the heterogeneous computing capabilities and communication links, these computation operations can be deployed onto different components in ISEA, resulting in various paradigms as detailed in Section V-B.
@@ V-B5 Hierarchical ISEA
The five ISEA paradigms—split, AI-on-server, AI-on-sensor, hierarchical, and cooperative—are not mutually exclusive but represent distinct characteristic dimensions of ISEA deployment. For instance, the first three paradigms (split, AI-on-server, and AI-on-sensor) define model placement strategies with trade-offs in hardware, communication, and computational complexity, while cooperative ISEA addresses multi-device coordination challenges. Crucially, these paradigms can be combined adaptively: hierarchical structures may employ different placement strategies across layers, and cooperative systems can integrate hybrid architectures (e.g., split for latency-sensitive tasks and AI-on-sensor for privacy constraints). Such flexibility, however, necessitates advanced scheduling algorithms and communication optimization, underscoring the need for context-aware ISEA design.
V-C Lessons Learned
  • The general architecture of ISEA includes sensors, edge brain, central brain and actuators. Sensors transform physical world information into digital signals. It is worth pointing out that sensors are not limited to standalone devices (e.g., cameras), but may coexist with other functions in shared hardware, such as RF sensing in ISAC systems. Edge brain encompasses all intelligent nodes in an edge network, which constantly communicate and collaborate with each other to accomplish ISEA tasks. The central brain, located in cloud data centers, focuses on time-insensitive operations such as global information aggregation and periodic model update distribution due to high access latency. The decisions made by edge brain are sent to actuators to interact with the physical environment.
  • Computation operations are integral to the execution of an ISEA task and can be deployed to various ISEA components based on device capabilities and collaboration mode. The characteristics of these computing operations vary substantially with the sensing data modality and AI backbone used[118]. As these operations are distributed across different nodes, communication interface between ISEA components are necessitated, including vertical data transmission for offloading or data aggregation and horizontal data exchange for peer collaboration.
  • ISEA instances can be categorized into different paradigms depending on the placement of AI models onto nodes, network topology and collaboration protocols. It is importance to notice that these paradigms can be combined or switched between in a shared ISEA system depending on the task requirements.
TABLE II: Summary of related works on digital air interface for ISEA
@@ V-C Lessons Learned
[136] A collaborative ISEA system where the BS aggregates data from multiple RF sensors Sensing disciminant gain Joint allocation of communication time, quantization bit and transmission energy
[137] A collaborative ISEA system where an access point (AP) aggregates data from multiple sensors and broadcasts the fused data Total execution time Information extractor design and a task-oriented channel allocation mechanism
@@
[142] Collaborative perception in vehicle networks E2E latency, perception probability Joint optimization of compression ratio, collaborator selection, offloading decision and computation resource allocation with Lyapunov-aided deep reinforcement learning
@@ VI-C1 Power Control and Bandwidth Allocation
The allocation of power and bandwidth resources to a given set of sensors shall consider the heterogeneous channels, data qualities, and computation capabilities of different sensors to meet the ISEA task performance requirement. In [134], an edge AI-assisted autonomous driving system is considered, where connected vehicles are scheduled to upload sensing data to multiple servers for perception model training. To maximize a defined metric, i.e., quality of training, the power and bandwidth allocation to different vehicles and time slots are jointly optimized. In [135], the authors consider ISAC settings where a BS simultaneously senses and transmits data packets, arriving in a Poisson process, to multiple users with a multi-functional OFDM signal. The subcarrier and power allocation problem is formulated to minimize the long-term average power consumption subject to sensing performance and queueing stability conditions and is solved via Lyapunov optimization. For multi-device cooperative RF sensing, the authors of [136] studies joint control of quantization bits, communication time allocation, and energy allocation to maximize the discriminant gain of sensing. In a collaborative ISEA system with a fusion center considered in[137], the inadequacy of uplink communication resources for uploading all sensing information necessitates information extraction, agent selection, and subchannel allocation according to channel states and importance of messages. The information extraction module is trained using the information bottleneck principle, the subchannel allocation is optimized to maximize the expected information gain, and message fusion is designed with a double-attention mechanism. The authors of [138] first propose a primal-descent dual ascent training algorithm for FL under hybrid data partition, which involves several groups of sensors with different types, to resolve the data dependence issue. Then, the optimization problem of sensor scheduling, bandwidth allocation, and quantization is formulated to minimize the training residual term and resource consumption, which is solved in a decentralized manner.
@@ VI-C3 Joint Computation-Communication Resource Allocation
In ISEA systems, computation and communication resources can both be bottlenecks of task performance, especially in view of time-varying channel conditions and limited on-sensor computation capability. This requires a joint controller coordinating communication and computation operations to ensure full resource utilization. In [140], a point-to-point split inference system involving a UAV and a BS is considered, where the UAV can switch between two modes, i.e., uploading raw data or extracted features, with different C2 resource consumption. Such mode switching is jointly optimized with the transmit power allocation for multiple timeslots and computational speed to minimize the total energy consumption. The resultant optimization problem is solved by embedding the optimal power allocation into a tiny reinforcement learning algorithm for optimal mode selection. In light of the tradeoff between reducing communication cost and preventing semantic loss as regulated by the compression ratio, in [141] a joint ratio selection and resource allocation problem is formulated for a multi-user semantic communication system aiming at supporting inference tasks for users. The solution approach is through alternative minimization that iterates between compression ratios and resource allocation. Further, a user selection algorithm is designed via branch and bound under communication resource constraints, considering that local tasks have different priority levels. Considering split ISEA where multi-user performs inference offloading with heterogeneous accuracy and latency requirements, the authors in [62] propose integrated batching and early exiting to maximize the system inference throughput. Under a total bandwidth constraint and different channel states, the optimized allocation of computation and communication resources for inference throughput maximization is realized via a low-complexity sub-optimal algorithm and a tree search-based optimal algorithm. For collaborative ISEA, the splitting point between sensor and server sub-models can be jointly optimized with C2 resources for minimized energy consumption subject to delay constraints[25]. For collaborative vehicles, [142] jointly optimizes collaborator selection, compression ratio, offloading decision, and computation resource allocation to minimize the E2E delay subject to a perception probability requirement.
@@ VII AirComp-based Air Interface for Integrated Sensing and Edge AI
TABLE III: Summary of Related Works on AirComp-based Air Interface for ISEA
@@
[166] Over-the-air FL under Byzantine attacks Convergence rate Implememt one-bit gradient quantization and majority vote to enhance robustness against Byzantine attackers
@@ VII-C1 Robust AirComp
In addition, AirComp performance is prone to interference from jammers or neighboring cells, which may lead to sudden drops in the sensing task performance, i.e., AI outage. A set of RRM techniques has been developed for interference management. In[163], a scheme called simultaneous signal-and-interference alignment is proposed to manage inference in a two-cell system. It divides the MIMO channel space into an interference alignment space and signal alignment space, and by constraining the interference into the former, the number of interference-free dimensions for AirComp can be maximized. In [145], the centralized power control problem for a multi-cell AirComp network is tackled by profiling the Pareto front of multi-cell MSEs and solving a sequence of second-order cone programming problems, and a distributed algorithm designed by exploiting the interference temperature technique. For multicluster networks, [164] developed a centralized branch-and-bound algorithm and a distributed algorithm for the transceiver optimization for transceiver design to minimize the aggregation MSE accounting for both signal misalignment error and interference. A novel technique, termed spectrum breathing, is developed in [165], which features cascaded gradient pruning and spread spectrum to reduce interference under a given bandwidth budget. In addition, in [166], one-bit gradient quantization and majority-vote aggregation have been proposed in over-the-air FL systems to counter Byzantine attacks.
@@ VIII Advanced Signal Processing for Integrated Sensing and Edge AI
TABLE IV: Summary of Related Works on Advanced Signal Processing for ISEA
@@
[208] Multi-user and multi-target ISAC Sum rate, sensing error Utilize constructive interference to combat multi-user interference and design ISAC waveforms to achieve the optimal Pareto front
@@ VIII-B ISEA with Multi-Functional Waveforms
From a physical layer perspective, advancing ISEA critically relies on the foundational design of waveforms. Multi-functional waveforms are sophisticated signal designs that integrate multiple functions, such as sensing, communication, localization, and radar, into a unified structure, leveraging advanced signal processing to optimize performance. Previous attempts in this field focused on extending the functionalities of existing systems. For example, communication symbols were modulated onto traditional radar waveforms, such as space-time coding over orthogonal radar waveforms, code shift keying using optimized coded radar signals, and embedding communication symbols in frequency-hopping multi-input multi-output (FH-MIMO) radar systems [198, 199, 200]. Some early works in this field explored the performance of sensing using signals generated through linear frequency modulation (LFM) and continuous phase modulation [201, 202]. These schemes aimed to utilize existing architectures to meet the limited demands of additional functionalities. More recent research focuses on the novel design of multi-functional waveforms for simultaneous monostatic sensing and communication (SMSC). In SMSC, a transceiver array emits dual-functional signals to illuminate the scene and gather environmental information through echo signals. At the same time, forward signals are reflected to a communication receiver for information delivery. The advantage of SMSC is that it allows the uncertainty of communication messages to be incorporated into the transmitted waveforms without compromising the coherent processing of monostatic sensing. This eliminates the need for cooperation overhead [203]. However, the distinct requirements of sensing and communication pose challenges in the design of multi-function waveforms, such as the Cramér-Rao bound (CRB) and data rates, leading to a tradeoff in performance [203, 221]. To address these tradeoffs, research has focused on beamforming techniques, where beam patterns are formed in a joint radar-communication MIMO system based on the directions of sensing targets and communication users to mitigate mutual interference [16, 205, 206, 207]. In multi-user and multi-target scenarios, multi-user interference is a critical challenge, which is addressed in [208] by introducing constructive interference. Additionally, time-frequency waveforms have been revisited to optimize subcarrier powers in a specific time-frequency region of interest, achieving a favorable performance tradeoff between radar and communications [209].
@@ IX-A3 Reconfigurable Intelligent Surfaces
  • Joint beamforming techniques: Reflection/refraction control of the RIS must be jointly designed with conventional transmit beamforming to maximize sensing performance, rather than simply focusing on maximizing SNR as in typical communication scenarios. For instance, [243] demonstrates how adaptive RIS beamforming can maximize the received signal power at RIS sensors—a crucial metric for sensing. Moreover, via controlled refraction/reflection of sensing signals, RIS allow receiving and distinguishing echo signals from multiple targets[244]. In general, beamforming and RIS control strategies should be explored to balance communication efficiency and sensing accuracy, as RIS allows manipulation of both communication and sensing signals[245].
@@ X-C Ultra-Low-Latency ISEA
However, existing SPT approaches targeting URLLC fall short of meeting the ultra-LoLa ISEA requirements. This shortcoming arises because ISEA applications are typically data-intensive and often require the transmission of high-dimensional features, which conflicts with SPT’s limited payload sizes and leads to communication bottlenecks. The problem is further exacerbated for tasks subject to strict latency requirements. Moreover, the E2E performance of sensing tasks cannot be accurately measured through decoding error probability alone, as that metric is more suited to assessing communication reliability. To address these challenges, the first framework that enables ultra-LoLa ISEA leverages the interplay between SPT and multi-view sensing—improving accuracy while adhering to stringent latency constraints [272]. Within such a framework, a fundamental tradeoff is identified and optimized between communication reliability and the number of sensing views, both of which are governed by the packet length. This integration of SPT and ISEA opens numerous opportunities for future work. One key direction is the simultaneous improvement of accuracy and reduction of latency through a joint optimization of packet length and coding rate, which requires further exploration into the effects of feature dimensionality and quantization on packet selection [273]. From an energy-efficiency standpoint, designing power control techniques that accommodate both fading channels and feature importance is also a promising avenue. Additionally, extending the ultra-LoLa inference paradigm to incorporate AirComp constitutes another compelling research direction.
@@
X-D Practical Challenges in ISEA
This subsection discusses challenges in practical implementation of ISEA from the perspectives of scalability, energy resource constraints, heterogeneous and dynamic networks and controlling overhead, highlighting key issues and potential research directions.
X-D1 Scalability
The scalability of ISEA systems is a critical challenge due to the massive number of sensors, edge devices, and AI models that must operate cohesively in 6G networks. As envisioned, 6G will support billions of interconnected devices, generating vast amounts of multi-modal sensory data (e.g., from RGB cameras, LiDARs, and RF sensors) that require real-time processing and transmission. This scale introduces several issues:
  • Cooperation of numerous devices: As more nodes in edge networks are equipped with AI and sensing capabilities, the number of sensors participating in cooperative ISEA is envisioned to continuously grow. For example, in cooperative autonomous driving scenarios, the number of agents in popular datasets have increased from $2$ in F-Cooper[275] to $6$ in OPV2V[276] to $12$ in V2X-Sim[277]. While such a trend enables more holistic sensing, it also increases communication overhead for multiple access and complicates scheduling and resource management. Moreover, coordinating distributed learning (e.g., federated learning) across thousands of devices introduces synchronization and convergence issues.
  • Data volume and processing overload: Developments in resolutions and ranges of sensing technologies leads to higher dimensionality of sensory data. For example, point clouds from LiDARs now have data rates up to several Gbps [35]. Sensors and edge brain with limited computational capabilities may struggle to process these data streams in real time. Also, the communication rate of data exchange grows proportionally, leading to a communication bottleneck.
  • Model scalability: State-of-the-art multi-modal foundation models consistently grows in the numbers of parameters and required computation operations, known as the scaling law. Deploying such models requires significant memory and computational resources, which may exceed the capabilities of resource-constrained edge devices and edge servers.
    Research directions: AirComp can be leveraged to aggregate sensory data efficiently [150], as its latency does not grow with the number of devices. However, with numerous devices, several challenges requires research efforts, such as device scheduling and synchronization. Neural network-based data down-sampling can be leverage to cope with the high data rate, and its interplay with wireless communication warrants future research. Model compression techniques, including quantization and pruning, can reduce the footprint of AI models, enabling their deployment on resource-limited devices[278]. Efficient inference techniques, such as speculative decoding and expert parallelism, can accelerate inference on large-scale models. Finally, hierarchical edge-cloud architectures can offload computationally intensive tasks to higher-tier nodes, balancing scalability and performance.
X-D2 Energy Constraints
Energy efficiency is an important concern for ISEA, given the resource-constrained nature of edge devices and sensors. The integration of sensing, AI inference, and communication in 6G networks demands significant energy, particularly for real-time, mission-critical applications.
  • Sensing energy costs: High-resolution sensors like LiDARs and cameras consume substantial power for data acquisition. For example, a typical LiDAR module can consume from 10 to 30 watts[279], posing challenges for battery-powered devices like drones or IoT sensors.
  • AI inference and training: Edge AI operations, including inference and distributed learning, are energy-intensive. Deep learning models, even when optimized, require significant computational resources for forward and backward passes, especially for real-time tasks like robotic control or semantic segmentation. Federated learning, while reducing data transmission, still incurs high energy costs for local model updates.
  • Communication energy costs: Transmitting high-dimensional sensory data or model updates over wireless channels consumes considerable energy, particularly in mmWave or THz bands, which require high transmit power to overcome path loss. The frequent data exchanges in ISEA for cooperative sensing or computation offloading further exacerbate energy demands.
    Research directions: Energy-efficient designs are critical for ISEA. Selective sensor wake-up based on query or environmental information can reduce the long-term average sensing power consumption, thereby extending the network lifetime. Energy-efficient JSCC can minimize transmission energy by allocating power across different messages [280] or adopt lightweight models to reduce the computation energy[281]. Dynamic power control for AirComp can optimize energy use during data aggregation [153]. Additionally, energy harvesting technologies, such as RF energy harvesting or solar-powered sensors, can extend the operational lifetime of devices. Finally, task offloading frameworks that balance local and edge computation based on energy availability can enhance overall efficiency.
X-D3 Heterogeneous and Dynamic Networks
The heterogeneity of 6G networks, encompassing diverse devices, sensing modalities, and communication protocols, poses significant interoperability challenges for ISEA. Seamless integration and cooperation among these components are essential for achieving optimal E2E task performance.
  • Diverse sensing modalities: ISEA integrates multiple sensing modalities (e.g., RF, LiDAR, and RGB-D cameras), each with distinct data formats, sampling rates, and processing requirements. Harmonizing these modalities for joint processing or fusion is challenging, especially when devices operate on different temporal and spatial scales. For instance, aligning RF sensing data and video streams with different streaming rates and updating frequencies requires synchronization mechanisms.
  • Heterogeneous devices and protocols: 6G networks will include a mix of legacy and next-generation devices, ranging from low-power IoT sensors to high-end autonomous vehicles. These devices may use different communication protocols (e.g., NB-IoT, LTE-M, 5G NR) and AI model formats, complicating cooperative tasks such as distributed inference or sensor data sharing. Moreover, hierarchical ISEA requires joint control and optimization across edge, transport and core networks.
  • Network dynamics: The dynamic nature of 6G networks, with high device mobility and varying channel conditions, affects sensing performance. For example, in vehicular networks, rapidly changing topologies due to high-speed vehicle movement can disrupt cooperative sensing or model synchronization, leading to degraded task performance.
    Research directions: To enhance interoperability, standardized data formats and protocols for multi-modal sensory data exchange are needed. Semantic communication can be used to reconcile different sensing modalities under a unified framework. Adaptive modulation and coding schemes tailored to sensing tasks can accommodate diverse device capabilities. Additionally, LLM-empowered network agents can dynamically manage resources and protocols to maintain seamless cooperation under different task requirements and network status[282].
X-D4 Controlling Overhead
Although the joint ISEA design enhances task performance, it also introduces additional overhead for signaling and computation operations in control algorithms. These overhead becomes particularly significant in ultra-low-latency or massive access applications.
  • Signaling overhead: The time-varying nature of sensing context requires frequent transmission of control decisions and feedback information between sensors and edge nodes. Meanwhile, the raised dynamic resource mapping between communication, AI and sensing introduces higher multiple access control overhead compared to 5G’s static scheduling[283]. In addition, while the unified waveform design in ISAC reduces redundant signal processing delays, it requires additional reference signals for cross-functional channel estimation (e.g., simultaneous CSI and radar sensing parameter extraction).
  • Computation overhead: Real-time cross-domain optimization demands more baseband processing due to 1) the coupled sensing-communication beamforming (e.g., solving high-dimensional Pareto fronts); 2) AI task QoS monitoring (e.g., perceptual quality metrics for generative AI). This places heavy computing burden on 6G network controllers.
    Research directions: Digital twins of the environment can reduce the information feedback to the minimum amount by incrementally updating the digital twin at the server. For example, the server can estimate the wireless channel using ray-tracing in the digital twin, which only requires low-dimensional object position feedback but not CSI feedback. The computation overhead of control algorithms can be simplified by approximated optimizations and decentralized computation. Moreover, advanced baseband hardware, e.g., in-memory computing[284], can increase processing speed by up to 1000$\times$ compared with existing semiconductor devices, making it a promising solution for ultra-low-latency applications.